OpenStudy (anonymous):

Who can help me with this question: A window washer stands on a scaffolding 30m above the ground.If he did 23,520J of work to reach the scaffolding,What is his mass?

OpenStudy (anonymous):

Hey sheelo, what level class is this for?

OpenStudy (anonymous):

I was wondering mainly the scope to explain this in, but while I have time I'll just do it the simple way. Since we're dealing strictly in the vertical direction (height in meters) the amount of work (W) to get the man to his current height is equal to his potential energy while standing at that height. As you may know potential energy (U) is equal to mass(m) times gravity (g) times height (h). With these facts we can setup an equation, with the units given in brackets. \[W = U [J] = m [kg] g \left[ \frac{ m }{ \sec ^{2} } \right]h[m]\]One thing to note here is that 1 Joule is equal to (kg * meters squared) / (second squared) or...\[1[J] = 1\frac{ [kg][m ^{2}] }{ [s ^{2}] }\]With your given initial work being in joules and the height in meters we just have to solve the equation for m and plug in our values while double checking that the units cancel.\[m = \frac{ U }{ gh } = \frac{23,520 \left[ \frac{ [kg][m ^{2} ]}{ s ^{2} } \right] }{9.81\left[ \frac{ m }{ s^{2} } \right]30\left[ m \right] }\]Everything except kg cancels, leaving us with an answer of around 80 kilograms. This would be roughly 175lbs which is a perfect reasonable weight for a man! Hope this helps.